There is 3 questions I have no idea how to solve. I want to know the steps if anyone can help.

1) How many mg of medication would you need to prepare 3000 ml of 1:2500 solution?

2) How many ml of 5% solution are needed to prepare 10 ml if 1:50 solution?

3)

Ordered: meidcatid t.i.d for 33 lb child

Label: medication 15mg/kg/day

How many mg can the child receive in one day?

How many mg can the child recieve in each dose?

Thank you in advance!
Answer #1

1) Molarity = Number of moles of solute/Volume of solutio in L

1/2500 = Number of moles of solute/3

Number of moles of solute = 3/2500 = 0.0012 moles

Multiply the moles by molar mass to get the amount in gms

2)

The number of moles must be same

M1V1 = M2V2

(5%)(V1) = 10(1/50)

V1 = 10 * 2/5 = 4 mL

Hence the volume required will be equal to 4 mL

3) 1 lbs = 0.453592 Kg

Child receive in one day = 15 mg/Kg/day * 33 lb * 0.453592 Kg/lb = 224.52 mg

Assuming it is given in the form of three dose, mg in each dose = 224.52 mg/3 = 74.85 mg